CYCLIC PRESENTATIONS OF THE TRIVIAL GROUP

MARTIN EDJVET, PAUL HAMMOND, AND NATHAN THOMAS

Abstract. We report on a computational group theory experiment involving a search for cyclic presentations of the trivial group. The list of such presentations obtained includes counterexamples to a conjecture of M J Dunwoody.